# VBIM-Net: Variational Born Iterative Network for Inverse Scattering Problems

### v1

@@ Abstract
Recently, studies have shown the potential of integrating field-type iterative methods with deep learning (DL) techniques in solving inverse scattering problems (ISPs). In this article, we propose a novel Variational Born Iterative Network, namely, VBIM-Net, to solve the full-wave ISPs with significantly improved flexibility and inversion quality. The proposed VBIM-Net emulates the alternating updates of the total electric field and the contrast in the variational Born iterative method (VBIM) by multiple layers of subnetworks. We embed the calculation of the contrast variation into each of the subnetworks, converting the scattered field residual into an approximate contrast variation and then enhancing it by a U-Net, thus avoiding the requirement of matched measurement dimension and grid resolution as in existing approaches. The total field and contrast of each layer’s output is supervised in the loss function of VBIM-Net, which guarantees the physical interpretability of variables of the subnetworks. In addition, we design a training scheme with extra noise to enhance the model’s stability. Extensive numerical results on synthetic and experimental data both verify the inversion quality, generalization ability, and robustness of the proposed VBIM-Net. This work may provide some new inspiration for the design of efficient field-type DL schemes.
@@ I Introduction
The second type of DL-based algorithm integrates traditional iterative methods with DNNs to introduce physical information into network design. Sanghvi et al. [32] designed CS-Net to learn the noise space component of the contrast source current and embedded it into two-fold SOM [33], which has a satisfying generalization ability but is still time-consuming for the iterative process. Liu et al. [34] transformed the ISP into a constrained optimization problem and designed PM-Net to emulate the alternating direction method of multiplier (ADMM) iterative process, updating the contrast source and contrast alternately. Zhang et al. [35] embedded CSI method into a convolutional neural network (CNN) structure. Liu et al. [36] proposed SOM-Net by replacing the contrast source update step in SOM iteration with a U-Net. In general, the aforementioned schemes all combine source-type iterative methods with DNN design, which achieve better generalization than typical image-to-image translation schemes. In parallel, there has been a few works exploring the potential of integrating DNNs with field-type iterative methods, treating the contrast and the total field as inversion variables. Beerappa et al. [37] embedded two cascaded CNNs into the iteration of DBIM, which reduces the reconstruction error of DBIM, but still demands heavy computation within iterations. In [38], Shan et al. proposed the neural Born iterative method (NeuralBIM) by utilizing the physics-informed supervised residual learning (PhiSRL) [39] to fully unroll the BIM iteration. However, NeuralBIM directly concatenates the scattering field residual and the contrast image as the network input, which in general requires the measurement dimension to well match the resolution of the inversion grid. More importantly, although NeuralBIM attempts to use multiple layers of subnetworks of the same structure to unroll the iterations, the parameters within each of the intermediate layers are not fully physically constrained. As a result, its data flow fails to well conform to that of the original iterations, thus degrading the model’s physical interpretability and generalization ability.
@@
In this paper, we propose a novel variational Born iterative network, namely, VBIM-Net, by unrolling the VBIM algorithm to solve the full-wave ISPs, aiming at overcoming the limitations of the existing field-type DL schemes in terms of efficiency, flexibility, and inverion quality. The proposed VBIM-Net emulates the alternating updates of the total electric field and the contrast in VBIM by multiple layers of subnetworks, avoiding the complexity of traditional iterative calculations. In each subnetwork, we transform the scattered field residual into the approximate contrast variation in VBIM, which naturally lifts the limitation on measurement dimension and gird resolution in NeuralBIM without sacrificing computational efficiency. The predicted contrast and total field of each subnetwork are explicitly supervised in the loss function to make the data flow conform to an iterative process, thereby enhancing the model’s reliability and interpretability. In the training stage, we adopt a hybrid dataset and design a training scheme with extra noise further to ensure the generalization ability and robustness of the model. Synthetic and experimental experiments both verify the superior performance of the proposed VBIM-Net over the existing ones. The main contributions of this work are summarized as follows.
@@
    The proposed VBIM-Net transforms the scattering field residual into approximate contrast variation through analytical computation. It aligns the input variables of the contrast update network in spatial dimensions and unifies them into the contrast domain. This avoids the requirement for matched measurement dimension and grid resolution in NeuralBIM and aids the network learning.
@@
    A loss function with a layer-wise constraint is designed, ensuring the outputs of each VBIM-Net layer to be consistent with the actual physical quantities, while providing more accurate physical quantity residuals for the next layer. This makes the data flow of VBIM-Net closer to a iterative algorithm, ensuring the reliability of the model.
@@
    In the training stage, we mix two synthetic datasets to enable the model to learn the inversion of homogeneous and inhomogeneous scatters. Besides, we implement a training scheme with extra noise and employ a loss function that scales with the signal-to-noise ratio (SNR), thereby enhancing the robustness of the model.
@@
    Extensive experiments have been performed on synthetic and experimental datasets, including challenging profiles with complex shapes, high contrast, and high noise level. The proposed VBIM-Net demonstrates significant advantages in terms of reconstruction quality, generalization ability, and robustness.
@@ II-C Born Iterative Method and Its Variational Version
which can be reduced to a LS estimation in noisy conditions,
$\boldsymbol{\chi}_{(k)}=\arg\min_{\boldsymbol{\chi}}{\left\|\mathbf{E}^{s}-\mathbf{G}_{S}\operatorname{diag}(\boldsymbol{\chi})\mathbf{E}^{t}_{(k)}\right\|^{2}}.$ (14)
In comparison, VBIM adopts the variational form of Eq. (13) in the $\boldsymbol{\chi}$-update step,
$\delta\mathbf{E}^{s}=\mathbf{G}_{S}\operatorname{diag}(\delta\boldsymbol{\chi})\mathbf{E}^{t}_{(k)},$ (15)
where $\delta\mathbf{E}^{s}=\mathbf{E}^{s}-\mathbf{E}^{s}_{(k)}$ is the residual of scattered field, $\mathbf{E}^{s}_{(k)}$ is the computed scattered field in the $k$th iteration,
@@ III The Proposed Variational Born Iterative Network
To fully leverage the nonlinear fitting capability of DNNs and avoid the computational overhead from iterative calculations, the proposed VBIM-Net simulates the VBIM iterations by multiple layers of subnetworks of the same structure. Each subnetwork embeds the contrast variation in VBIM to transfer the information from scattered field residual. In order to make the variables in each subnetwork correspond to physical quantities, we introduce a layer-wise constraint in the loss function, which ensures the model’s reliability. Furthermore, we employ a training scheme with extra noise to enhance the robustness of the model.
@@ III-A Network Architecture
As depicted in Fig. 2, VBIM-Net includes $K$ layers of subnetworks to simulate the iterative process of VBIM. Assume that the initial value of contrast $\boldsymbol{\chi}_{(0)}$ is calculated by BPS, and the initial value of total field $\mathbf{E}^{t}_{(0)}$ is set according to the first-order Born approximation [12], i.e., $\mathbf{E}^{t}_{(0)}=\mathbf{E}^{i}$. Here, we concatenate the incident field $\mathbf{E}^{i}$ of $N_{i}$ incidences to obtain a real-value tensor with $2N_{i}$ channels as the input to VBIM-Net, and the network outputs the estimated total field for $N_{i}$ incidences at each layer. This allows the network to fully couple the data from multiple incidences to reconstruct the contrast of unknown scatterers. For lossless scatters, the channel number $C$ of $\boldsymbol{\chi}$ is 1, while for lossy scatters, $C=2$.
Each layer of VBIM-Net consists of a $\mathbf{E}^{t}$-update step and a $\boldsymbol{\chi}$-update step. In the $\mathbf{E}^{t}$-update step, instead of using the full-wave forward calculation in Eq. (12), we adopt the form of PhiSRL [39], which emulates the fixed-point iterative method by a neural network (NN) with an input-output residual connection. In the $k$th layer of the VBIM-Net, the $\mathbf{E}^{t}$-update step first calculates the incident field residual
@@
In the $\boldsymbol{\chi}$-update step, we first estimate the variation $\delta\boldsymbol{\chi}$ of the contrast according to Eq. (15). By stacking the data of $N_{i}$ incidences, we can transform Eq. (15) into the following linear equation,
@@
where $(\cdot)^{+}$ denotes the ReLU function.
The structure of each layer of VBIM-Net is shown in Fig. 3. We adopt U-Net [26] to build NNs $\mathcal{F}^{\delta\mathbf{E}^{t}}_{k}$ and $\mathcal{F}^{\delta\boldsymbol{\chi}}_{k}$ for predicing the variation $\delta{\mathbf{E}^{t}_{(k)}}$ and $\delta{\boldsymbol{\chi}_{(k)}}$, which efficiently integrates multi-level feature information and leverages the similarity between input and output, and has been widely applied to solve nonlinear ISPs [28], [30], [31], [40]. The spatial dimension and channel number of each layer of U-Net are marked in Fig. 3, where the number of input channel and output channel are denoted by $C_{\mathrm{in}}$ and $C_{\mathrm{out}}$, respectively. For the total field update network $\mathcal{F}^{\delta\mathbf{E}^{t}}_{k}$, $C_{\mathrm{in}}=4N_{i},C_{\mathrm{out}}=2N_{i}$. For the contrast update network $\mathcal{F}^{\delta\boldsymbol{\chi}}_{k}$, $C_{\mathrm{in}}=2,C_{\mathrm{out}}=1$ for lossless scatterers, and $C_{\mathrm{in}}=4,C_{\mathrm{out}}=2$ for lossy scatterers. The CNNs in the U-Net consist of $3\times 3$ convolutions, batch normalization (BN), and ReLU activation functions, and the final output layers use $1\times 1$ convolution to project the feature map to the output space. Algorithm 1 summarizes the data flow of VBIM-Net. In this paper, VBIM-Net is assumed to consist of 7 layers of subnetworks, which share the same structure but have independent parameters.
@@ III-B Loss Function
In order to make each layer output of VBIM-Net corresponds to the actual physical quantity, we designed layer-wise constraint in the loss function, which explicitly constrain the output contrast and total field of each layer. The overall loss function is formulated as
@@
with constant $c\in[0,1)$, which is set to 0.8 by default. In this way, the outputs of each subnetwork in VBIM-Net correspond to actual physical quantities, while also providing reasonable residual $\delta\mathbf{E}^{i}$ and $\delta\mathbf{E}^{s}$ for the next subnetwork. In general, the role of the layer-wise constraint is a kind of regularization, which assists the data flow of VBIM-Net to consistent with the paradigm of iterative algorithms. In the following section, we will validate the importance of the layer-wise constraint.
@@ IV Numerical Results
In this section, we verify the effectiveness of VBIM-Net on synthetic and experimental data, and evaluate the model’s reconstruction quality, generalization ability, and robustness.
Figure 5: Examples of lossy scatterers in the synthetic dataset.
(a) Cylinder dataset example.
(b) MNIST dataset example.
@@ IV-A Benchmark Algorithms and Performance Metrics
We select SOM [20] and NeuralBIM [38] as benchmark algorithms. SOM is a classic source-type iterative method that decomposes the contrast current into a deterministic part and an ambiguous part through SVD, and uses a two-step CG algorithm to alternately optimize the contrast and ambiguous current component, which has good inversion performance in most scenarios [1], [20], [36].
@@ IV-B System Configurations and Training Details
Parameters (VBIM-Net) 29.24 Million
Parameters (modified NeuralBIM) 48.62 Million
@@
TABLE II: Performance Metrics of Reconstruction Results for Test#1-Test#10
@@
The system configurations of the synthetic data is as follows: $N_{i}$ line sources and $N_{r}$ line receivers are equally placed on a circle with a radius of 1.67 m, where $N_{i}=16$, $N_{r}=32$ by default, and the other measurement dimensions will be discussed in Section IV-D. The frequency of the incident wave is 3 GHz. $D$ is the DOI of size $0.2\mathrm{m}\times 0.2\mathrm{m}$ with a free space background. In the forward problem, $D$ is discretized into a grid of $100\times 100$, and the MoM is used to simulate the scattered field $\mathbf{E}^{s}$, while in the inverse problem, $D$ is discretized into a grid of $64\times 64$ to avoid the inverse crime. The simulated $\mathbf{E}^{s}$ is noise-free, so we add different levels of Gaussian noise to $\mathbf{E}^{s}$ to simulate the actual measurement data. The noise level is defined as $(\|\mathbf{N}\|_{F}/\|\mathbf{E}^{s}\|_{F})$, where $\mathbf{N}$ is the noise matrix. During training, we randomly add 5%-35% noise to $\mathbf{E}^{s}$ and employ the weighted loss in Eq. (33). By default, we assume that the scatterers are lossless, and we will discuss the inversion results for lossy scatterers separately. The system configuration for the experimental data is slightly different, which will be described in Section IV-E.

### v2

@@ Abstract
Recently, studies have shown the potential of integrating field-type iterative methods with deep learning (DL) techniques in solving inverse scattering problems (ISPs). In this article, we propose a novel Variational Born Iterative Network, namely, VBIM-Net, to solve the full-wave ISPs with significantly improved structural rationality and inversion quality. The proposed VBIM-Net emulates the alternating updates of the total electric field and the contrast in the variational Born iterative method (VBIM) by multiple layers of subnetworks. We embed the analytical calculation of the contrast variation into each subnetwork, converting the scattered field residual into an approximate contrast variation and then enhancing it by a U-Net, thus avoiding the requirement of matched measurement dimension and grid resolution as in existing approaches. The total field and contrast of each layer’s output is supervised in the loss function of VBIM-Net, imposing soft physical constraints on the variables in the subnetworks, which benefits the model’s performance. In addition, we design a training scheme with extra noise to enhance the model’s stability. Extensive numerical results on synthetic and experimental data both verify the inversion quality, generalization ability, and robustness of the proposed VBIM-Net. This work may provide some new inspiration for the design of efficient field-type DL schemes.
@@ I Introduction
The second type of DL-based algorithm integrates traditional iterative methods with DNNs to introduce physical information into network design. Sanghvi et al. [32] designed CS-Net to learn the noise space component of the contrast source current and embedded it into two-fold SOM [33], which has a satisfying generalization ability but is still time-consuming for the iterative process. Liu et al. [34] transformed the ISP into a constrained optimization problem and designed PM-Net to emulate the alternating direction method of multiplier (ADMM) iterative process, updating the contrast source and contrast alternately. Zhang et al. [35] embedded CSI method into a convolutional neural network (CNN) structure. Liu et al. [36] proposed SOM-Net by replacing the contrast source update step in SOM iteration with a U-Net. In general, the aforementioned schemes all combine source-type iterative methods with DNN design, which achieve better generalization than typical image-to-image translation schemes. In parallel, there has been a few works exploring the potential of integrating DNNs with field-type iterative methods, treating the contrast and the total field as inversion variables. Beerappa et al. [37] embedded two cascaded CNNs into the iteration of DBIM, which reduces the reconstruction error of DBIM, but still demands heavy computation within iterations. In [38], Shan et al. proposed the neural Born iterative method (NeuralBIM) by utilizing the physics-informed supervised residual learning (PhiSRL) [39] to fully unroll the BIM iteration. However, NeuralBIM directly concatenates the scattered field residual and the contrast image as the network input, which in general requires the measurement dimension to well match the resolution of the inversion grid. More importantly, although NeuralBIM attempts to use multiple layers of subnetworks with the same structure to emulate the iterations, the computations and outputs of each intermediate layer are not governed by any form of physical constraints. As a result, its data flow fails to well conform to that of the original iterations, thus degrading the model’s physical interpretability and generalization ability.
In this paper, we propose a novel variational Born iterative network, namely, VBIM-Net, by unrolling the VBIM algorithm to solve the full-wave ISPs.
@@
We embed physics-inspired analytical computations into the network architecture and incorporate soft physical constraints into the loss function, aiming to overcome the drawbacks of the existing field-type DL schemes in terms of structural rationality and inversion quality. In each subnetwork, we transform the scattered field residual into an approximate contrast variation inspired by VBIM, which naturally lifts the limitation on measurement dimension and gird resolution in NeuralBIM without sacrificing computational efficiency. Through the proposed layer-wise constraint, the predicted contrast and total field in each subnetwork are explicitly supervised in the loss function, guiding the data flow to conform to an iterative process and thereby enhancing the model’s reliability and interpretability. In the training stage, we adopt a hybrid dataset and design a training scheme with extra noise further to ensure the generalization ability and robustness of the model. Synthetic and experimental experiments both verify the superior performance of the proposed VBIM-Net over the existing ones. The main contributions of this work are summarized as follows.
@@
    The proposed VBIM-Net transforms the scattered field residual into approximate contrast variation through analytical computation. It aligns the input variables of the contrast update network in spatial dimensions and unifies them into the contrast domain. This avoids the requirement for matched measurement dimension and grid resolution in NeuralBIM, while fully utilizing the information from the scattered field residual to update the contrast, which improves the structural rationality of the existing field-type DL scheme.
@@
    A loss function with a layer-wise constraint is designed, ensuring the outputs of each VBIM-Net layer to be consistent with the actual physical quantities, while providing more accurate physical quantity residuals for the next layer. This makes the data flow of VBIM-Net closer to a iterative algorithm, thereby ensuring the reliability and generalization performance when training the deep VBIM-Net model.
@@
    In the training stage, we mix two synthetic datasets to enable the model to learn the inversion of homogeneous and inhomogeneous scatterers. Besides, we implement a training scheme with extra noise and employ a loss function that scales with the signal-to-noise ratio (SNR), thereby enhancing the robustness of the model.
@@
    Extensive experiments have been performed on synthetic and experimental datasets, including challenging profiles with complex shapes, high contrast, and high noise level. The proposed VBIM-Net demonstrates significant advantages in terms of reconstruction quality, generalization ability, and robustness. We also conduct a series of experiments to verify the effectiveness and necessity of the novel designs proposed in VBIM-Net.
@@ II-C Born Iterative Method and Its Variational Version
which can be reduced to a LS estimation in noisy conditions. To alleviate the ill-posedness of ISPs, L2 regularization is commonly introduced, which can be formulated as
$\boldsymbol{\chi}_{(k)}=\arg\min_{\boldsymbol{\chi}}{\left\{\|\mathbf{E}^{s}-\mathbf{G}_{S}\operatorname{diag}(\boldsymbol{\chi})\mathbf{E}^{t}_{(k)}\|_{2}^{2}+\lambda\cdot\|\boldsymbol{\chi}\|_{2}^{2}\right\}}.$ (14)
As an evolved version of BIM, VBIM adopts the variational form of Eq. (13) in the $\boldsymbol{\chi}$-update step,
$\delta\mathbf{E}^{s}_{(k)}=\mathbf{G}_{S}\operatorname{diag}(\delta\boldsymbol{\chi})\mathbf{E}^{t}_{(k)},$ (15)
where $\delta\mathbf{E}^{s}_{(k)}=\mathbf{E}^{s}-\mathbf{E}^{s}_{(k)}$ is the residual of scattered field, $\mathbf{E}^{s}_{(k)}$ is the computed scattered field in the $k$th iteration,
@@ III The Proposed Variational Born Iterative Network
To fully leverage the nonlinear fitting capability of DNNs and avoid the computational overhead from iterative calculations, the proposed VBIM-Net simulates the VBIM iterations by multiple layers of subnetworks with the same structure. Each subnetwork embeds the contrast variation in VBIM to transfer the information from scattered field residual. In order to make the variables in each subnetwork correspond to physical quantities, we introduce a layer-wise constraint in the loss function, which ensures the model’s reliability. Furthermore, we employ a training scheme with extra noise to enhance the robustness of the model.
@@ III-A Network Architecture
As depicted in Fig. 2, VBIM-Net includes $K$ layers of subnetworks to simulate the iterative process of VBIM. Assume that the initial value of contrast $\boldsymbol{\chi}_{(0)}$ is calculated by BPS, and the initial value of total field $\mathbf{E}^{t}_{(0)}$ is set according to the first-order Born approximation [12], i.e., $\mathbf{E}^{t}_{(0)}=\mathbf{E}^{i}$. Here, we concatenate the incident field $\mathbf{E}^{i}$ of $N_{i}$ incidences to obtain a real-value tensor with $2N_{i}$ channels as the input to VBIM-Net, and the network outputs the estimated total field for $N_{i}$ incidences at each layer. This allows the network to fully couple the data from multiple incidences to reconstruct the contrast of unknown scatterers. For contrast $\boldsymbol{\chi}$, we represent it with images with dimension $C\times M\times M$. For lossless scatters, the channel number $C$ is 1, while for lossy scatters, $C=2$.
The overall architecture of VBIM-Net refers to PhiSRL in [39]. To solve the matrix equation $\mathbf{Ax}=\mathbf{b}$, PhiSRL stacks multiple layers of neural networks with the same structure and residual connections to simulate the fixed-point iteration update process. In the $k$-th layer, the neural network computes
$\mathbf{x}_{(k)}=\mathbf{x}_{(k-1)}+\mathcal{F}_{k}(\delta_{(k-1)},\Theta_{k})=\mathbf{x}_{(k-1)}+\delta\mathbf{x}_{(k)},$ (17)
where $\delta_{(k-1)}=\mathbf{b}-\mathbf{A}\mathbf{x}_{(k-1)}$ is the equation residual from the $(k-1)$-th layer, $\mathcal{F}_{k}$ is the neural network in the $k$-th layer, and $\delta\mathbf{x}_{(k)}=\mathcal{F}_{k}(\delta_{(k-1)},\Theta_{k})$ is the predicted variation for $\mathbf{x}$.
Each layer of VBIM-Net consists of a $\mathbf{E}^{t}$-update step and a $\boldsymbol{\chi}$-update step. In the $\mathbf{E}^{t}$-update step, instead of using the full-wave forward calculation in Eq. (12), we adopt the form of one-step PhiSRL. Specifically, in the $k$th layer of the VBIM-Net, the $\mathbf{E}^{t}$-update step first calculates the incident field residual
@@
In the $\boldsymbol{\chi}$-update step, we also draw on the idea of PhiSRL, but based on the VBIM algorithm, we convert the residual $\delta{\mathbf{E}^{s}_{(k)}}$, which originally belongs to the field domain, into the contrast domain. Specifically, we first estimate the variation $\delta\boldsymbol{\chi}$ of the contrast according to Eq. (15). By stacking the data of $N_{i}$ incidences, we can transform Eq. (15) into the following linear equation,
@@
where $(\cdot)^{+}$ denotes the ReLU function. Given the convergence advantage of VBIM over BIM, our proposed approximate contrast variation more efficiently conveys the information needed for contrast updating than the scattered field residuals in NeuralBIM, while naturally avoiding the $N_{r}\times N_{i}=M\times M$ constraint.
The structure of each layer of VBIM-Net is shown in Fig. 3. We adopt U-Net [26] to build NNs $\mathcal{F}^{\delta\mathbf{E}^{t}}_{k}$ and $\mathcal{F}^{\delta\boldsymbol{\chi}}_{k}$ for predicting the variation $\delta{\mathbf{E}^{t}_{(k)}}$ and $\delta{\boldsymbol{\chi}_{(k)}}$, which efficiently integrates multi-level feature information and leverages the similarity between input and output, and has been widely applied to solve nonlinear ISPs [28], [30], [31], [40]. The spatial dimension and channel number of each layer of U-Net are marked in Fig. 3, where the number of input channel and output channel are denoted by $C_{\mathrm{in}}$ and $C_{\mathrm{out}}$, respectively. For the total field update network $\mathcal{F}^{\delta\mathbf{E}^{t}}_{k}$, $C_{\mathrm{in}}=4N_{i},C_{\mathrm{out}}=2N_{i}$. For the contrast update network $\mathcal{F}^{\delta\boldsymbol{\chi}}_{k}$, $C_{\mathrm{in}}=2,C_{\mathrm{out}}=1$ for lossless scatterers, and $C_{\mathrm{in}}=4,C_{\mathrm{out}}=2$ for lossy scatterers. The CNNs in the U-Net consist of $3\times 3$ convolutions, batch normalization (BN), and ReLU activation functions, and the final output layers use $1\times 1$ convolution to project the feature map to the output space. Algorithm 1 summarizes the data flow of VBIM-Net. In this paper, VBIM-Net is assumed to consist of 7 layers of subnetworks, which share the same structure but have independent parameters.
@@ III-B Loss Function
In order to make each layer output of VBIM-Net corresponds to the actual physical quantity, we designed a layer-wise constraint in the loss function, which explicitly constrains the output contrast and total field of each layer, decaying the importance of each layer’s output from back to front. The overall loss function is formulated as
@@
with parameter $c\in[0,1)$ as a decay factor, which is set to 0.8 by default. In this way, the outputs of each subnetwork in VBIM-Net correspond to actual physical quantities, while also providing reasonable residual $\delta\mathbf{E}^{i}$ and $\delta\mathbf{E}^{s}$ for the next subnetwork. In general, the layer-wise constraint can be interpreted as a regularization technique based on soft physical constraints, which assists the data flow of VBIM-Net to consistent with the paradigm of iterative algorithms. In the following section, we will verify the effectiveness and necessity of the layer-wise constraint in training deep VBIM-Net models.
@@ III-C Complexity Analysis
Figure 5: Examples of scatterers in the synthetic dataset for lossy scatterers. The first row also represents the examples from the lossless dataset.
@@ IV Numerical Results
In this section, we demonstrate the effectiveness of VBIM-Net on synthetic and experimental data, evaluating the model’s reconstruction quality, generalization ability, and robustness. Based on this, we designed a series of experiments to validate the physical insight and experimental efficacy of the new designs in proposed VBIM-Net.
@@ IV-A Benchmark Algorithms and Performance Metrics
We select BIM, VBIM, SOM [20] and NeuralBIM [38] as benchmark algorithms. The BIM and VBIM algorithm are introduced in Section II-C, and the L2 regularization parameter $\lambda$ is empirically set to $5\times 10^{-4}$ here.
SOM is a classic source-type iterative method that decomposes the contrast current into a deterministic part and an ambiguous part through SVD, and uses a two-step CG algorithm to alternately optimize the contrast and ambiguous current component, which has good inversion performance in most scenarios [1], [20], [36].
@@
(a) Cylinder dataset example.
(b) MNIST dataset example.
@@ IV-B System Configurations and Training Details
The system configurations of the synthetic data are as follows: $N_{i}$ line sources and $N_{r}$ line receivers are equally placed on a circle with a radius of 1.67 m, where $N_{i}=16$, $N_{r}=32$ by default, and the other measurement dimensions will be discussed in Section IV-D. The frequency of the incident wave is 3 GHz. $D$ is the DOI of size $0.2\mathrm{m}\times 0.2\mathrm{m}$ with a free space background. In the forward problem, $D$ is discretized into a grid of $100\times 100$, and the MoM is used to simulate the scattered field $\mathbf{E}^{s}$. In the inverse problem, we need to discretize $D$ at a lower resolution than the forward problem to avoid the inverse crime[1], and $64\times 64$ is chosen in this paper. The simulated $\mathbf{E}^{s}$ is noise-free, so we add different levels of Gaussian noise to $\mathbf{E}^{s}$ to simulate the actual measurement data. The noise level is defined as $(\|\mathbf{N}\|_{F}/\|\mathbf{E}^{s}\|_{F})$, where $\mathbf{N}$ is the noise matrix. During training, we randomly add 5%-35% noise to $\mathbf{E}^{s}$ and employ the weighted loss in Eq. (34). By default, we assume that the scatterers are lossless, and we will discuss the inversion results for lossy scatterers separately in Section IV-D. The system and dataset configuration for the experimental data is slightly different, which will be described in Section IV-E.
We mix the MNIST [42] and cylinder dataset at a ratio of 1:1 to obtain the synthetic dataset, in order to increase the richness of the data, making the network to learn inversion tasks for homogeneous and inhomogeneous scatterers with various shapes. For the cylinder dataset, each scene contains 1 to 3 cylinders with a 0.01m-0.05m radius. For lossless scatters, the contrast range is $[0.2,2.2]$. For lossy scatters, the range of the real part of the contrast is $[0.2,2.2]$, and the range of the imaginary part is $[0,1]$. Fig. 5 shows several examples from the lossy dataset, with the first row in the figure also representing examples from the lossless dataset. Unless otherwise specified, in the subsequent experiments, we use the lossless dataset for training and testing by default. We randomly generate 20,000 data samples, of which 80% are for training and 20% for testing.
VBIM-Net and the modified NeuralBIM are implemented in PyTorch and computed on a Nvidia V100 GPU, and they share the same training configuration. Table I summarizes the implementation details of both, and Table II lists the number of parameters, total training time and single-sample inference time for both. Note that since the modified NeuralBIM adds a trainable dimension transformation module to implement the mapping $\delta\mathbf{E}^{s}\rightarrow\mathbf{F}^{\delta\mathbf{E}^{s}}$, the number of trainable parameters of the modified NeuralBIM is larger than that of VBIM-Net, which also leads to a longer training time. However, during the inference stage, the $\boldsymbol{\chi}$-update network of VBIM-Net contains an analytical computation of the approximate contrast variation $\delta\tilde{\boldsymbol{\chi}}$, resulting in a longer inference time.
@@
TABLE II: The number of parameters, total training time and single-sample inference time for modified NeuralBIM and VBIM-Net
Method Parameters Total training time Single-sample inference time
NeuralBIM 48.62 Million 16h 0.0178s
VBIM-Net 29.24 Million 10.58h 0.0273s
@@
TABLE III: Performance Metrics of Reconstruction Results for Test#1-Test#10
@@
IV-C Synthetic Data Inversion: Within-Dataset Test
In this subsection, we first visually display the output of each VBIM-Net layer, and then test the model’s performance on the MNIST and cylinder datasets. Finally, we conduct a series of experiments to verify the efficacy of the subnetwork input configuration and the layer-wise constraint in VBIM-Net.
IV-C1 Updated Contrast Reconstructions of VBIM-Net
IV-C2 Within-Dataset Samples Reconstruction
Fig. 7 shows several reconstruction results on the cylinder and MNIST test sets of these three algorithms when the noise level is 10%, and the corresponding performance metrics are summarized in Table III. Reconstruction results indicate that VBIM-Net outperforms all the comparison algorithms. Although both VBIM-Net and NeuralBIM are designed based on the field-type iterative method, the reconstruction quality of VBIM-Net is significantly better. This is mainly attributed to the analytical calculation of the approximate contrast variation $\delta\tilde{\boldsymbol{\chi}}$ that converts the scattered field residual $\delta\mathbf{E}^{s}$ into the contrast domain, making it more conducive to feature extraction by the subsequent CNN. Moreover, the U-Net used by VBIM-Net is more effective in extracting the multi-level features in contrast images than the ResNet in NeuralBIM, which also helps to improve the model performance.
IV-C3 Validation of the Efficacy of the Subnetwork Input Configuration
To verify the effectiveness of the subnetwork input configuration in VBIM-Net, we trained the following four models with different subnetwork inputs while keeping all other training conditions consistent:
  1. 1.
    The original VBIM-Net, where $\delta\tilde{\boldsymbol{\chi}}_{(k)}\oplus\boldsymbol{\chi}_{(k-1)}$ and $\delta\mathbf{E}^{i}_{(k)}\oplus\mathbf{E}^{t}_{(k-1)}$ are used as inputs for the $k$-th $\boldsymbol{\chi}$-update and $\mathbf{E}^{t}$-update networks, respectively;
  2. 2.
    Removing the input $\boldsymbol{\chi}_{(k-1)}$ from the $\boldsymbol{\chi}$-update network of VBIM-Net;
  3. 3.
    Removing the input $\mathbf{E}^{t}_{(k-1)}$ from the $\mathbf{E}^{t}$-update network of VBIM-Net;
  4. 4.
    Removing both $\boldsymbol{\chi}_{(k-1)}$ and $\mathbf{E}^{t}_{(k-1)}$, keeping only $\delta\tilde{\boldsymbol{\chi}}_{(k)}$ and $\delta\mathbf{E}^{i}_{(k)}$ as the inputs for the $\boldsymbol{\chi}$-update and $\mathbf{E}^{t}$-update networks, respectively.
The SSIM performance of these four models on the test set are shown in Fig. 9, and the first (Q1) and third (Q3) quartiles are marked. It can be seen that the original VBIM-Net performs the best, while removing $\boldsymbol{\chi}_{(k-1)}$ or $\mathbf{E}^{t}_{(k-1)}$ leads to a loss in performance and stability. This proves that in VBIM-Net, using the physical quantities output from the previous layer as inputs to the current layer helps the model better learn the updates of these quantities.
Furthermore, we find that removing $\boldsymbol{\chi}_{(k-1)}$, meaning that only $\delta\tilde{\boldsymbol{\chi}}_{(k)}$ is used as the input for the $\boldsymbol{\chi}$-update network, has a relatively minor impact on model performance, which indicates that the approximate contrast variation inspired by VBIM provides sufficient information for VBIM-Net to compute the contrast updates, demonstrating VBIM-Net’s advantage in structural rationality.
Figure 9: The SSIM performance on the test set of VBIM-Net after removing certain input physical quantities. The quartiles Q1 and Q3 are marked.
IV-C4 Validation of the Efficacy of the Layer-Wise Constraint
In order to verify the efficacy of the layer-wise constraint, we perform the following ablation experiment: set $c=0$ in Eq. (30) to only supervise the output of the last layer of VBIM-Net (where $10^{0}=$ 1 is defined), while keeping other settings unchanged. Fig. 10 shows the contrast MSE on the test set during training with/without the layer-wise constraint. It also shows the training convergence curve of the final layer output under $c=0.8$. When using the layer-wise constraint, the convergence trends of the training set and the test set are essentially consistent, indicating minimal signs of overfitting. Additionally, in this case, the reconstruction quality can progressively improve with the stacking of update blocks. With the same number of subnetwork layers, the model with the layer-wise constraint can converge to a better performance. Fig. 11 illustrates the reconstruction results without the layer-wise constraint, where the scatterer is the same as in Fig. 6(a). It can be seen that removing the layer-wise constraint can result in an ambiguous physical interpretation of the output from intermediate layers, as well as a deterioration in the final reconstruction quality.